OpenStudy (anonymous):

One day a store sold  26  sweatshirts. White ones cost  $9.95  and yellow ones cost  $11.50.   In all,  $275.75  worth of sweatshirts were sold. How many of each color were sold?

Directrix (directrix):

Let W represent a white SS Let Y represent a yellow SS

Directrix (directrix):

White ones cost $9.95 so 9.95*W dollars worth of white ss were sold Yellow ones cost $11.50 so 11.50*Y dollars worth of yellow ss were sold

OpenStudy (anonymous):

9.95w + 11.50y = 275.75 is how I would set it up?

Directrix (directrix):

Yes if you want to work this problem in two equations in two variables. If you want to work the problem in only one equation, then no.

OpenStudy (anonymous):

so which method would I use? the elimination or sub?

Directrix (directrix):

9.95W + 11.50Y = 275.75 W + Y = 26

Directrix (directrix):

Sub would probably be easier. It's your call.

OpenStudy (anonymous):

I suck at that method... elimination is my choice

Directrix (directrix):

9.95W + 11.50Y = 275.75 W + Y = 26 ---> W = 26 - Y 9.95* (26 - Y) + 11.50Y = 275.75

Directrix (directrix):

9.95* (26 - Y) + 11.50Y = 275.75 ------------------------ Okay, you can do elimination.

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

how would I solve it using one variable?

Directrix (directrix):

You said you would solve the problem by elimination. I said okay. I had already started with substitution. So, I substituted and then used an online calculator to solve for Y. So, if you want to stick with substitution, I'll go back and do that method by hand.

Directrix (directrix):

One Variable Approach: 9.95* (26 - Y) + 11.50Y = 275.75 ---> Solve that.

OpenStudy (anonymous):

lets just do it your way because I am confused

Directrix (directrix):

9.95W + 11.50Y = 275.75 W + Y = 26 --------------> ---> W = 26 - Y ------------ In the first equation, substitute (26 - Y) for W 9.95* (26 - Y) + 11.50Y = 275.75 258.70 - 9.95Y + 11.50Y = 275.75

OpenStudy (anonymous):

258.7

Directrix (directrix):

- 9.95Y + 11.50Y = 275.75 -258.70 1.55*Y = 17.05 Y = ? What is 17.05 divided by 1.55 ? @dontre

OpenStudy (anonymous):

11

Directrix (directrix):

Y = 11 so 11 yellow shirt. How many white ones?

Directrix (directrix):

W + Y = 26 ---> W = 26 - Y

Directrix (directrix):

@dontre Please answer the question so that we can get a wrap.

OpenStudy (anonymous):

is it 15?

Directrix (directrix):

Correct.

Directrix (directrix):

11 Yellows and 15 Whites